# Quantum chaos and the arrow of time

### v8

@@ Abstract
Summary. We show that quantum chaotic systems inherently possess an arrow of time. Working at leading order in perturbation theory, we show that the second law of thermodynamics constrains the off-diagonal terms in ETH statement. The new constraint has to be taken into account when one is working beyond probe limit. It would be worthwhile to study the higher order terms in perturbation theory and see how they constrain higher point functions.
Acknowledgements.
@@
The author would like to thank Gautam Mandal for helpful discussions during one of which the main idea for this paper arose. The author is fully supported by the Department of Science and Technology (Government of India) under the INSPIRE Faculty fellowship scheme (Ref. No. DST/INSPIRE/04/2020/002105).

### v9

@@ 6 Conclusions
We show that quantum chaotic systems inherently possess an arrow of time. Working at leading order in perturbation theory, we show that the second law of thermodynamics constrains the off-diagonal terms in ETH statement. The new constraint has to be taken into account when one is working beyond probe limit. If the system is at a given energy $E$, the energy eigenstates which are at a distance greater than $W/2$ in the energy spectrum play a very important role in the chaotic dynamics of the system. It has already been shown that these states constrain the Lyapunov exponent [11].
We also calculated the higher order terms of $\lambda(t)$ in $\Delta E$. For reasonable perturbation strength, the higher order terms are small compare to the leading term. The odd power terms of $\lambda(t)$ are identically zero.
We also study perturbations with large values of $\lambda(t)$ numerically. We considered perturbations of the form $\lambda(t)=\lambda\delta(t)$ and $\lambda(t)=2\;\text{rect}(t/2d)$ which is the rectangular function of height $2$ and width $2d$. $\Delta E$ and $\Delta S$ as a function of $\lambda$ and $d$ are plotted in Figure 9. $\Delta E$ remains always positive (negative) for positive (negative) initial temperature. This agrees with the Kelvin’s form of the second law of thermodynamics. The change in entropy is also always positive in these examples. Note that the energy even jumps the infinite temperature region.
Figure 9: Left: Change in energy after large perturbations in the chaotic XXZ spin chain. Starting from the positive temperature state, the system always gains energy (red curves). Starting from the negative temperature state, the system always loses energy (blue curves). Right: The change in entropy is always postive.
We can still reduced the entropy by using a fine-tuned perturbation. This can be accomplished by predominantly turning on modes $\tilde{\lambda}(\omega)$ which allows transition to only lower energy levels. But this fine-tuning requires the knowledge of a large part of the spectrum and precise knowledge of the initial state. This is reminiscent of the Maxwell’s demon problem in classical thermodynamics. In the present case, finding the energy levels at the required precision involves performing highly precise measurements which increases the total entropy of the system and the outside agent.
Acknowledgements
@@
The author would like to thank Gautam Mandal for helpful discussions during one of which the main idea for this paper arose.
Funding information
The author is fully supported by the Department of Science and Technology (Government of India) under the INSPIRE Faculty fellowship scheme
(Ref. No. DST/INSPIRE/04/2020/002105).